OpenStudy (anonymous):

find intercepts of equation y=2x - 8

OpenStudy (anonymous):

Its in slope intercept form y=mx+b where be is the y intercept so the y intercept is (0, -8) to find the x intercept, plug in 0 for y and then solve for x

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

so would it be (4,0)

OpenStudy (anonymous):

(4,0),(0,-8)

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

can you help with this?

OpenStudy (anonymous):

write an equation in point slope form of the line that passes through the given point and with the given slope m

OpenStudy (anonymous):

(-6,2) m=3

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) where m is slope and (x1, y1) is your point

OpenStudy (anonymous):

y-2=3(x-5)

OpenStudy (anonymous):

Close, y-2=3(x+6)

OpenStudy (anonymous):

but it was a negative 6

OpenStudy (anonymous):

Since theres a negative in the formula and its a negative 6, it becomes positive.

OpenStudy (anonymous):

thanks! also one more.

OpenStudy (anonymous):

use point by point plotting to sketch the graph of the function

OpenStudy (anonymous):

f(x)=32/x

OpenStudy (anonymous):

can you just tell me how to do that? how to plug it in my calc

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, you should be able to just input the equation. instead of f(x)= it might say y= just put in y=32/x and it should graph it for you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome